# Cascaded Self-Evaluation Augmented Training for Efficient Multimodal Large Language Models

### v1

@@ 3.3 Cas-SEAT
4 Experiments
4.1 Experimental Setup
4.1.1 Datasets
We train on MathV360K dataset and evaluate on Math-Vista, Math-V, We-Math. MathV360K is a widely used dataset for MLLM training. Math-Vista), Math-V, We-Math are three widely used public benchmarks for MLLM evaluation.
MathVista
Method Average Multi-choice Free-form Text Integer General VQA Math-targeted VQA
Base (inference) 0.2850 0.4407 0.1022 0.4407 0.1124 0.3391 0.2389
@@ 4.1.1 Datasets
Finetune (inference) 0.3160 0.4481 0.1609 0.4481 0.1770 0.3196 0.3130
Finetune (evaluation) 0.3490 0.4926 0.1804 0.4926 0.1986 0.3543 0.3444
CoT (inference) 0.3380 0.4815 0.1696 0.4815 0.1866 0.3326 0.3426
[…]
SEAT (inference) 0.2760 0.4278 0.0978 0.4278 0.1077 0.2957 0.2593
SEAT (evaluation) 0.2850 0.4389 0.1043 0.4389 0.1148 0.3196 0.2556
Cas-SEAT (inference) 0.3390 0.4889 0.1630 0.4889 0.1794 0.3652 0.3167
Cas-SEAT (evaluation) 0.4500 0.6222 0.2478 0.6222 0.2727 0.4848 0.4204
Improve 19.68% 15.07% 31.04% 15.07% 31.04% 11.50% 17.01%
Table 1: Comparison of Cas-SEAT and the Baseline on MathVista Dataset
Math-V
Method All Level1 Level2 Level3 Level4 Level5 ALG ARI CG COM
Base (inference) 0.0526 0.0800 0.0690 0.0364 0.0444 0.0299 0.0000 0.0000 0.2941 0.0526
@@
SEAT (inference) 0.0592 0.1132 0.0488 0.0357 0.0000 0.0882 $0$ul0.0526 0.1579 0.0000
SEAT (evaluation) 0.0888 0.1509 0.0732 0.0714 0.0000 0.1324 $0$ul0.0526 0.1579 0.0526
Cas-SEAT (inference) 0.1711 0.1321 0.1341 0.2321 0.1778 0.1912 0.1579 0.1579 0.2632 0.1579
Cas-SEAT (evaluation) 0.2763 0.2642 0.2439 0.2500 0.3111 0.3235 0.3158 0.1579 0.4211 0.2105
Improve 55.57% 27.33% 25.01% 179.96% 40.01% 57.12% 199.91% 200.19% 14.31% 99.91%
Table 2: Comparison of Cas-SEAT and the Baseline on Math-V Dataset
@@
Method Avg(Strict) IG(Strict) CM(Strict) Avg(Loose) IG(Loose) CM(Loose)
Base (inference) 0.0143 0.0171 0.0057 0.0657 0.0171 0.0571
@@
Finetune (inference) 0.0695 0.0362 0.0514 0.2562 0.0362 0.2381
Finetune (evaluation) 0.0695 0.0362 0.0514 0.2562 0.0362 0.2381
CoT (inference) 0.0438 0.0305 0.0286 0.1752 0.0305 0.1600
@@
SEAT (inference) 0.0105 0.0095 0.0057 0.0429 0.0095 0.0381
SEAT (evaluation) 0.0114 0.0114 0.0057 0.0495 0.0114 0.0438
Cas-SEAT (inference) 0.0533 0.0343 0.0362 0.2114 0.0343 0.1943
Cas-SEAT (evaluation) 0.1076 0.0438 0.0857 0.3495 0.0438 0.3276
@@
Table 3: Comparison of Cas-SEAT and the baselines on We-Math dataset
4.1.2 Baselines
To verify the applicability, the following EMLLMs are implemented and compared with the counterparts combined with the proposed method. We primarily analyzed the effectiveness of our method based on the EMLLMs LLaVA-v1.5(7B) Liu et al. (2024), Qwen2-VL(2B) Wang et al. (2024). Since current EMLLMs research often involves training on various datasets, including those frequently used as test datasets, such as the ones we selected, and many MLLM studies do not disclose their training datasets, we opted not to include EMLLMs published after the release of these datasets to ensure a fair comparison.
4.1.3 Implementation Details
“Base”: The pretrained MLLM Liu et al. (2024); Wang et al. (2024). “Finetune”: The pretrained MLLM is lora fine-tuned on Math360k Hu et al. (2022). “CoT”: Qwen2-VL (7B) performs CoT reasoning on Math360k. The MLLM then learns from this data Wang et al. (2023); Wadhwa et al. (2024). “SEAT”: The pretrained MLLM is fine-tuned on Math360k and subsequently performs reasoning on Math360k. Qwen2-VL (7B) evaluates the reasoning outputs of the MLLM, and the MLLM learns from this evaluated data Xiong et al. (2024). “Cas-SEAT”: The pretrained MLLM is fine-tuned on Math360k and subsequently performs reasoning on Math360k. Qwen2-VL (7B) evaluates the MLLM’s incorrect reasoning outputs. These evaluated data are combined with CoT data, and the MLLM learns from this combined dataset.
@@ 4.1.3 Implementation Details
Method UCU AL CPF UPF CSF USF BTF One-step Two-step Three-step
@@
[…]
SEAT (inference) 0.0942 0.0263 0.1099 0.1476 0.1083 0.1624 0.1176 0.1210 0.1361 0.1091
SEAT (evaluation) 0.0942 0.0263 0.1179 0.1476 0.1130 0.1624 0.1176 0.1243 0.1472 0.1091
Cas-SEAT (inference) 0.2679 0.3439 0.3240 0.3216 0.3230 0.2408 0.3240 0.3103 0.2861 0.2364
Cas-SEAT (evaluation) 0.4256 0.1982 0.4294 0.3546 0.3303 0.3431 0.3462 0.3909 0.3528 0.4788
Improve 8.32% -14.42% 24.36% 16.76% 11.44% 4.60% 34.71% 20.54% 13.40% 27.41%
Table 4: Detailed comparison of Cas-SEAT and the baselines on We-Math dataset
@@
After obtaining outputs on the test dataset, we first use a more powerful MLLM to extract answers for the MathVista dataset, followed by using regular expressions to calculate accuracy. For Math-V and We-Math, we directly use the regular expressions provided in the datasets to calculate accuracy.
4.2 Experimental Results
4.2.1 Overall Assessment of Cas-SEAT
As shown in Tables 1, 2, 3, and 4, we analyzed the experimental results on the MathVista, Math-V, and We-Math datasets. We evaluated the accuracy of two types of outputs for each method: direct inference (inference) and self-evaluation (evaluation). In addition to overall performance, we conducted tests on multiple subsets of the three datasets, including generalization capability, data types, problem difficulty, and task types. Below is the correspondence between these categories and the columns in the tables. (Due to dataset limitations, some categories could not be tested for certain datasets.): Overall Performance: MathVista (Table 1: Average), Math-V (Table 2: All), We-Math (Table 3: Score (Strict), Score (Loose)). Cas-SEAT shows very significant improvements over the baseline on all datasets, achieving around a 50% improvement on both Math-V and We-Math. Generalization Capability: We-Math (Table 3: Inadequate Generalization (Strict), Complete Mastery (Strict), Inadequate Generalization (Loose), Complete Mastery (Loose)). The improvement range of 9.5% to 55.25% demonstrates Cas-SEAT’s strong generalization capabilities, excelling at problems involving more integrated knowledge points, as evidenced by the significant improvement in Complete Mastery (CM). Data Types: MathVista (Table 1: Text, Integer). Cas-SEAT is better at answering integer-type questions. Problem Difficulty: MathVista (Table 1: Multi-choice, Free-form), Math-V (Table 2: Level1, Level2, Level3, Level4, Level5), We-Math (Table 3: One-step, Two-step, Three-step). The Math-V subsets are divided into five difficulty levels, with higher levels indicating more complex problems. The We-Math subsets are classified according to the number of reasoning steps, with more steps indicating higher difficulty. Cas-SEAT shows a more pronounced advantage on more challenging problems. Task Types: MathVista: General VQA, Math-targeted VQA; Math-V: Algebra (ALG), Arithmetic (ARI), Combinatorial Geometry (CG), Combinatorics (COM); We-Math: Understanding and Conversion of Units (UCU), Angles and Length (AL), Calculation of Plane Figures (CPF), Understanding of Plane Figures (UPF), Calculation of Solid Figures (CSF), Understanding of Solid Figures (USF), Basic Transformations of Figures (BTF). Cas-SEAT exhibits significant improvements over the best baseline method on almost all problem types, especially for numerical computation tasks (Math-V: ALG, ARI). For VQA tasks, Cas-SEAT performs better on the more difficult Math-targeted VQA.
In summary, we found that: 1. After fine-tuning, the inference ability of EMLLMs improves, but their self-evaluation ability declines. This is because fine-tuning enhances their mathematical reasoning ability but leads to forgetting general knowledge. Using synthetic CoT data based on mathematical datasets for enhanced training shows a similar pattern. 2. SEAT, which integrates CoT inference and self-evaluation training, suffers a significant drop in CoT inference ability, without a substantial improvement in self-evaluation. Even if EMLLMs are first trained with synthetic CoT data, using only self-evaluation training in the second round significantly reduces their CoT reasoning ability. This demonstrates the detrimental effect of long prompts and outputs on EMLLMs, and the difficulty of maintaining both CoT inference and self-evaluation capabilities simultaneously. 3. Cas-SEAT, by separating the CoT inference and self-evaluation tasks and shortening the prompt, preserves robust inference ability. More importantly, it achieves far stronger self-evaluation performance than the baselines. Our approach Cas-SEAT achieves remarkably significant improvements over the best baseline methods, both overall and across each subset. The detailed analysis shows that Cas-SEAT is especially adept at more difficult numerical computation tasks.
4.2.2 Evaluation on Different EMLLMs
Figure 4: Comparison of Cas-SEAT and the Baseline based on Qwen2-VL(2B).
@@ 4.2.2 Evaluation on Different EMLLMs
Figure 5: A bar chart analysis sample in MathVista. Green background indicates the raw data, red text represents incorrect reasoning processes (sometimes with no reasoning process), pink background and yellow background denote results from direct reasoning and self-evaluation, respectively. Blue text and blue background indicate the corrected reasoning process and corrected results, respectively.
4.2.3 Case Study

### v2

@@ 3.4. Cas-SEAT
Table 1. Comparison of Cas-SEAT and the Baselines on MMMU Dataset. BUS, TE, AD, HM, SCI and HSS respectively denotes the Business, Tech and Engineering, Art and Design, Health and Medicine, Science, Humanities and Social Science.
MMMU
Model Extra Data Total Accuracy BUS TE AD HM SCI HSS
Heuristics Baselines
Random Choice Base 0.2210 0.2470 0.2140 0.2920 0.2070 0.1800 0.2000
Open-Source Models Inference
miniGPT4-7B Base 0.2680 0.2130 0.2380 0.2920 0.3070 0.2870 0.2920
CogVLM-17B Base 0.3210 0.2560 0.2890 0.3800 0.3120 0.2510 0.4150
LLaVA-v1.5-13B Base 0.3640 0.2270 0.3140 0.5170 0.3870 0.2930 0.5330
@@
Finetune 0.3559 0.2600 0.3571 0.5133 0.3067 0.2800 0.4333
CoT 0.3118 0.1800 0.2666 0.4200 0.2800 0.2733 0.4916
SEAT 0.3204 0.2400 0.2857 0.4066 0.2800 0.2600 0.5000
LLaVA-v1.5-7B Cas-SEAT 0.3226 0.2667 0.3238 0.3733 0.3400 0.2267 0.4250
Open-Source Models Self-Evaluation
@@
SEAT 0.3462 0.2600 0.3333 0.4133 0.3066 0.2866 0.5166
LLaVA-v1.5-7B Cas-SEAT 0.5193 0.4933 0.5000 0.6333 0.4533 0.4266 0.6416
Improve 31.60% 48.00% 36.35% 11.75% 15.26% 39.09% 22.21%
Table 2. Comparison of Cas-SEAT and the Baselines on MathVista Dataset. FQA, MWP, ALG, ARI, LOG, NUM and STA respectively denote figure QA, math word problem, algebraic, arithmetic, logical, numeric, and statistical.
MathVista
Model Extra Data Average FQA MWP ALG ARI LOG NUM STA
Close-Source Models Inference
Gemini 1.0 Nano 2 Base 0.3060 0.2860 0.3060 0.2710 0.2980 0.1080 0.2080 0.3350
Gemini 1.0 Pro Base 0.4520 0.4760 0.3920 0.4520 0.3880 0.1080 0.3260 0.5680
GPT-4V Base 0.4990 0.4310 0.5750 0.5300 0.4900 0.2160 0.2010 0.5580
Open-Source Models Inference
InstructBLIP-7B Base 0.2530 0.2310 0.1830 0.2180 0.2710 0.1890 0.2040 0.2310
LLaVA-13B Base 0.2610 0.2680 0.1610 0.2730 0.2010 0.2430 0.1830 0.2510
SPHINX-V1-13B Base 0.2750 0.2340 0.2150 0.2560 0.2810 0.1620 0.1740 0.2360
LLaVA-v1.5-13B Base 0.324 0.2677 0.2366 0.3701 0.272 0.1622 0.2639 0.2558
[…]
Finetune 0.3160 0.2416 0.3011 0.3665 0.2890 0.1081 0.2431 0.2724
CoT 0.3380 0.2416 0.2903 0.4413 0.2805 0.1081 0.2153 0.2658
SEAT 0.2760 0.1970 0.1667 0.3665 0.2181 0.0811 0.1667 0.2093
LLaVA-v1.5-7B Cas-SEAT 0.3390 0.3011 0.2581 0.3986 0.2805 0.1622 0.1806 0.3023
Open-Source Models Self-Evaluation
Base 0.3530 0.2974 0.1613 0.4021 0.3088 0.1622 0.2500 0.2857
@@
SEAT 0.3490 0.3048 0.2634 0.4413 0.2493 0.2162 0.1736 0.2990
LLaVA-v1.5-7B Cas-SEAT 0.4500 0.4201 0.4032 0.4733 0.4023 0.4054 0.2986 0.4086
Improve 19.68% 37.83% 24.98% 6.41% 13.61% 87.51% 7.49% 35.16%
4. Experiments
4.1. Experimental Setup
4.1.1. Datasets
We train on MathV360K (Shi et al., 2024) dataset and evaluate on MMMU (Yue et al., 2024), Math-Vista (Lu et al., 2023), Math-V (Wang et al., 2024b), We-Math (Qiao et al., 2024). MathV360K is a widely used dataset for MLLM training. MMMU is a widely used public benchmarks for MLLM evaluation and Math-Vista, Math-V, We-Math are three widely used public benchmarks for MLLM evaluation in math domain.
Table 3. Comparison of Cas-SEAT and the Baselines on We-Math dataset
@@ 4.1.1. Datasets
Model Extra Data Avg(Strict) IG(Strict) CM(Strict) Avg(Loose) IG(Loose) CM(Loose)
Close-Source Models Inference
GPT-4V Base 0.3105 0.1448 0.2381 0.5143 0.1448 0.0333
Gemini 1.5 Pro Base 0.2638 0.1124 0.2076 0.4600 0.1124 0.1203
Qwen-VL-Max Base 0.1048 0.0762 0.0667 0.2552 0.0762 0.2028
Open-Source Models Inference
LLaVA-1.5-13B Base 0.0248 0.0114 0.019 0.0952 0.014 0.0895
LLaVA-v1.6-13B Base 0.0524 0.0324 0.0362 0.2200 0.0324 0.2621
DeepSeek-VL-7B Base 0.0629 0.0457 0.0400 0.2095 0.0457 0.2899
G-LLaVA-13B Base 0.0648 0.0457 0.0419 0.2229 0.0457 0.3598
@@
Finetune 0.0695 0.0362 0.0514 0.2562 0.0362 0.2381
CoT 0.0438 0.0305 0.0286 0.1752 0.0305 0.1600
SEAT 0.0105 0.0095 0.0057 0.0429 0.0095 0.0381
LLaVA-v1.5-7B Cas-SEAT 0.0533 0.0343 0.0362 0.2114 0.0343 0.1943
Open-Source Models Self-Evaluation
@@
Finetune 0.0695 0.0362 0.0514 0.2562 0.0362 0.2381
[…]
SEAT 0.0114 0.0114 0.0057 0.0495 0.0114 0.0438
LLaVA-v1.5-7B Cas-SEAT 0.1076 0.0438 0.0857 0.3495 0.0438 0.3276
@@
Table 4. Comparison of Cas-SEAT and the Baselines across multiple dimensions on the MathVista Dataset.
MathVista
Model Extra Data Multi-choice Free-form Text Integer General VQA Math-targeted VQA
Base 0.4407 0.1022 0.4407 0.1124 0.3391 0.2389
Finetune 0.4481 0.1609 0.4481 0.1770 0.3196 0.3130
CoT 0.4815 0.1696 0.4815 0.1866 0.3326 0.3426
SEAT 0.4278 0.0978 0.4278 0.1077 0.2957 0.2593
LLaVA-v1.5-7B Cas-SEAT 0.4889 0.1630 0.4889 0.1794 0.3652 0.3167
Open-Source Models Self-Evaluation
[…]
Finetune 0.4926 0.1804 0.4926 0.1986 0.3543 0.3444
[…]
SEAT 0.4389 0.1043 0.4389 0.1148 0.3196 0.2556
LLaVA-v1.5-7B Cas-SEAT 0.6222 0.2478 0.6222 0.2727 0.4848 0.4204
Improve 15.07% 31.04% 15.07% 31.04% 11.50% 17.01%
Table 5. Detailed comparison of Cas-SEAT and the Baselines on We-Math dataset. UCU, AL, CPF, UPF, CSF, USF, and BTF respectively denote Understanding and Conversion of Units, Angles and Length, Calculation of Plane Figures, Understanding of Plane Figures, Calculation of Solid Figures, Understanding of Solid Figures, and Basic Transformations of Figures.
@@
Model Extra Data UCU AL CPF UPF CSF USF BTF
Open-Source Models Inference
[…]
Finetune 0.3770 0.1649 0.3196 0.2664 0.2723 0.2371 0.1950
[…]
SEAT 0.0942 0.0263 0.1099 0.1476 0.1083 0.1624 0.1176
LLaVA-v1.5-7B Cas-SEAT 0.2679 0.3439 0.3240 0.3216 0.3230 0.2408 0.3240
Open-Source Models Self-Evaluation
[…]
Finetune 0.3770 0.1649 0.3196 0.2664 0.2723 0.2371 0.2172
[…]
SEAT 0.0942 0.0263 0.1179 0.1476 0.1130 0.1624 0.1176
LLaVA-v1.5-7B Cas-SEAT 0.4256 0.1982 0.4294 0.3546 0.3303 0.3431 0.3462
Improve 8.32% -14.42% 24.36% 16.76% 11.44% 4.60% 34.71%
Table 6. Comparison on tasks of varying difficulty.
We-Math Math-V
Model Extra Data One-step Two-step Three-step All Level1 Level2 Level3 Level4 Level5
Open-Source Models Inference
Base 0.1621 0.1472 0.1394 0.0526 0.0800 0.0690 0.0364 0.0444 0.0299
Finetune 0.3004 0.2750 0.3394 0.1743 0.2075 0.1951 0.0893 0.1778 0.1912
CoT 0.2905 0.2500 0.2000 0.1414 0.2075 0.1951 0.0893 0.1778 0.1912
SEAT 0.1210 0.1361 0.1091 0.0757 0.1400 0.0690 0.0364 0.0444 0.0896
LLaVA-v1.5-7B Cas-SEAT 0.3103 0.2861 0.2364 0.1711 0.2075 0.1951 0.0893 0.1778 0.2059
Open-Source Models Self-Evaluation
@@
SEAT 0.1243 0.1472 0.1091 0.1447 0.1509 0.1098 0.0714 0.2222 0.1912
LLaVA-v1.5-7B Cas-SEAT 0.3909 0.3528 0.4788 0.2763 0.2642 0.2439 0.2500 0.3111 0.3235
Improve 20.54% 13.40% 27.41% 55.57% 27.33% 25.01% 7.71% 40.01% 57.12%
4.1.2. Baselines
To verify the applicability, the following EMLLMs are implemented and compared with the counterparts combined with the proposed method. We primarily analyzed the effectiveness of our method based on the EMLLMs LLaVA-v1.5(7B) (Liu et al., 2024a), Qwen2-VL(2B) (Wang et al., 2024a). Since current EMLLMs research often involves training on various datasets, including those frequently used as test datasets, such as the ones we selected, and many MLLM studies do not disclose their training datasets, we opted not to include EMLLMs published after the release of these datasets to ensure a fair comparison.
We also included the following models as references: miniGPT4 (Zhu et al., 2023), CogVLM (Wang et al., 2023a), LLaVA-v1.6 (Liu et al., 2024b), Gemini 1.0 (Team et al., 2023), Gemini 1.5 (Team et al., 2023), GPT-4V (OpenAI, 2023), Shared GPT-4V (Chen et al., 2024), SPHINX-V1 (Lin et al., 2023), G-LLaVA (Gao et al., 2023), DeepSeek-VL (Lu et al., 2024), Qwen-VL (Bai et al., 2023).
4.1.3. Implementation Details
@@
“Base”: The pretrained MLLM (Liu et al., 2024a; Wang et al., 2024a). “Finetune”: The pretrained MLLM is lora fine-tuned on Math360k (Hu et al., 2022). “CoT”: Qwen2-VL (7B) performs CoT reasoning on Math360k. The MLLM then learns from this data (Wang et al., 2023b; Wadhwa et al., 2024). “SEAT”: The pretrained MLLM is fine-tuned on Math360k and subsequently performs reasoning on Math360k. Qwen2-VL (7B) evaluates the reasoning outputs of the MLLM, and the MLLM learns from this evaluated data (Xiong et al., 2024). “Cas-SEAT”: The pretrained MLLM is fine-tuned on Math360k and subsequently performs reasoning on Math360k. Qwen2-VL (7B) evaluates the MLLM’s incorrect reasoning outputs. These evaluated data are combined with CoT data, and the MLLM learns from this combined dataset. After obtaining outputs on the test dataset, we first use a more powerful MLLM to extract answers for the MathVista dataset, followed by using regular expressions to calculate accuracy. For Math-V and We-Math, we directly use the regular expressions provided in the datasets to calculate accuracy.
@@
Figure 4. A bar chart analysis sample in MathVista. Green background indicates the raw data, red text represents incorrect reasoning processes (sometimes with no reasoning process), pink background and yellow background denote results from direct reasoning and self-evaluation, respectively. Blue text and blue background indicate the corrected reasoning process and corrected results, respectively.
4.2. Experimental Results
4.2.1. Overall Assessment of Cas-SEAT
As shown in Tables 1, 2, 3, 4, 5, 6, we analyze the experimental results on the MMMU, MathVista, Math-V, and We-Math datasets. We evaluated the accuracy of two types of outputs for each method: direct inference (inference) and self-evaluation (evaluation). In addition to overall performance, we conducted tests on multiple subsets of these datasets, including generalization capability, data types, problem difficulty, and task types. Below is the correspondence between these categories and the columns in the tables. (Due to dataset limitations, some categories could not be tested for certain datasets.): Overall Performance: MMMU (Table 1: Total Accuracy), MathVista (Table 2: Average), We-Math (Table 3: Score (Strict), Score (Loose)). Cas-SEAT shows very significant improvements over the baseline, achieving around a 20% $\sim$ 50% improvement on both the aforementioned datasets, which validates the effectiveness of the proposed approach. Furthermore, we conducted a detailed comparison of the performance under different Data Types, Problem Difficulty, and Task Types. Data Types: MathVista (Table 4: Text, Integer). Cas-SEAT is better at answering integer-type questions. Problem Difficulty: As shown in Table 6, the Math-V subsets are divided into five difficulty levels, with higher levels indicating more complex problems. The We-Math subsets are classified according to the number of reasoning steps, with more steps indicating higher difficulty. Cas-SEAT shows a more pronounced advantage on more challenging problems. Task Types: MMMU (Table 1): BUS, TE, AD, HM, SCI, HSS; MathVista (Table LABEL:tab:main_mathvista,tab:main_mathvista_2): General VQA, Math-targeted VQA, FQA, MWP, ALG, ARI, LOG, NUM, STA; We-Math (Table 5): UCU, AL, CPF, UPF, CSF, USF, BTF. Math-V (Table 11 in Appendix): ALG, ARI, CG, COM; Cas-SEAT exhibits significant improvements over the best baseline method on almost all problem types, especially for numerical computation tasks (Math-V: ALG, ARI). For VQA tasks, Cas-SEAT performs better on the more difficult Math-targeted VQA.
In summary, we found that: (1). After fine-tuning, the inference ability of EMLLMs improves, but their self-evaluation ability declines. This is because fine-tuning enhances their mathematical reasoning ability but leads to forgetting general knowledge. Using synthetic CoT data based on mathematical datasets for enhanced training shows a similar pattern. (2). SEAT, which integrates CoT inference and self-evaluation training, suffers a significant drop in CoT inference ability, without a substantial improvement in self-evaluation. Even if EMLLMs are first trained with synthetic CoT data, using only self-evaluation training in the second round significantly reduces their CoT reasoning ability. This demonstrates the detrimental effect of long prompts and outputs on EMLLMs, and the difficulty of maintaining both CoT inference and self-evaluation capabilities simultaneously. (3). Cas-SEAT, by separating the CoT inference and self-evaluation tasks and shortening the prompt, preserves robust inference ability. More importantly, it achieves far stronger self-evaluation performance than the baselines. Our approach Cas-SEAT achieves remarkably significant improvements over the best baseline methods, both overall and across each subset. The detailed analysis shows that Cas-SEAT is especially adept at more difficult numerical computation tasks.
4.2.2. Case Study